OpenStudy (anonymous):

Matt plans to put concrete on a rectangular portion of his driveway. The portion is 8 feet long and 4 inches high . The price of concrete is $98 per cubic yard. The total cost of the concrete Matt needs is $58.07 . Which of the following is closest to the width of the portion of the driveway on which Matt plans to put concrete? [1 foot=12 inches; 1 yard = 3 feet] (1point)

OpenStudy (anonymous):

Can anyone HELP !

OpenStudy (ivettef365):

what are the choices ?

OpenStudy (anonymous):

1 yard=3 feet=36(3*12)inches 3 feet equal 3/3=1 yard 8 feet equal 8/3 yard 36 inches equal 36/36=1 yard 4 inches equal 4/36=1/9 yard so 8/3*1/9*width=58.07/98 so width = 2 yard

OpenStudy (anonymous):

Choices are - A.0.5feet B.1.5feet C.3feet D.6feet